OpenStudy (anonymous):

3 + 1 +1/3 +1/9 +. . diverge or converge

OpenStudy (anonymous):

ratio=1/3

OpenStudy (anonymous):

geometric series with common ratio of \(\frac{1}{3}\) and since \(\frac{1}{3}<1\) it converges